OpenStudy (anonymous):

Apples An apple farmer finds that he has to throw out 15 bad apples from the 225 he has picked. What is the experimental probability that the next apple he picks will be good?

OpenStudy (btaylor):

\[\frac{225-15}{225} = \frac{210}{225} = 93.33%\]

jimthompson5910 (jim_thompson5910):

The experimental probability of getting a bad apple is simply P(bad apple) = (# of bad apples)/(# of apples total) = 15/225 = 1/15 Now subtract that from 1 to get the answer 1 - 1/15 = 14/15